# A Novel Energy-Efficient Cross-Layer Design for Scheduling and Routing in 6TiSCH Networks

### v1

@@ Abstract
The 6TiSCH protocol stack was proposed to ensure high-performance communications in the Industrial Internet of Things (IIoT). However, the lack of sufficient time slots for nodes outside the 6TiSCH’s Destination Oriented Directed Acyclic Graph (DODAG) to transmit their Destination Advertisement Object (DAO) messages and cell reservation requests significantly hinders their integration into the DODAG. This oversight not only prolongs the device’s join time but also increases energy consumption during the network formation phase. Moreover, challenges emerge due to the substantial number of control packets employed by both the 6TiSCH Scheduling Function (SF) and routing protocol (RPL), thus draining more energy resources, increasing medium contention, and decreasing spatial reuse. Furthermore, an SF that overlooks previously allocated slots when assigning new ones to the same node may increase jitter, and more complications ensue when it neglects the state of the TSCH queue, thus leading to packet dropping due to queue saturation. Additional complexity arises when the RPL disregards the new parent’s schedule saturation during parent switching, which results in inefficient energy and time usage. To address these issues, we introduce in this paper novel mechanisms, strategically situated at the intersection of SF and RPL that are designed to balance the control packet distribution and adaptively manage parent switching. Our proposal, implemented within the 6TiSCH simulator, demonstrates significant improvements across vital performance metrics, such as node’s joining time, jitter, latency, energy consumption, and amount of traffic, in comparison to the conventional 6TiSCH benchmark.
Index Terms:
Industrial IoT, TSCH, 6TiSCH, SF, RPL network formation, scheduling, routing.
I Introduction
Amid current digital advancements, the presence of the Internet permeates nearly every aspect of our lives, spanning domains from smart cities and agriculture to industry and medicine. At the heart of this expansive reach is the Internet of Things (IoT), poised to interlink billions of devices with the traditional Internet. Many of these devices, while cost-effective, come with inherent constraints related to energy, computation, storage, and communication capabilities. The overarching success of IoT hangs on these resource-limited objects’ ability to operate reliably and sustainably while maintaining connectivity [1]. In IoT systems, energy efficiency stands at the forefront of technological challenges, with researchers globally striving to meld enhanced performances with reduced or equilibrated energy consumption. Specifically, the medium access control (MAC) layer is pivotal in controlling the energy drain since it ensures the network’s reliability by regulating access to the shared communication medium; typically a wireless channel. The IEEE.802.15.4e standard has been crafted to cater to industrial applications. It introduces different operational MAC modes, such as Time Slotted Channel Hopping (TSCH), Deterministic and Synchronous Multi-channel Extension (DSME), and Low Latency Deterministic Network (LLDN) [2], [3]. Among these modes, TSCH stands out as the leading one and is included in several standards such as ANSI/ISA100.11a and WirelessHart IEC62591 [4], [5].
@@ I Introduction
The core concept of TSCH lies in the integration of Time Division Multiple Access (TDMA) with channel hopping techniques. This fusion results in a dual medium access mechanism based on time and channel. In this context, “time” is referred to by time slot, and “frequency” is referred to by channel. A time slot within a specific channel is defined as a cell. For every time slot, shown in Fig. 1, an IoT node specifies its activity in a given channel, choosing to transmit to a node (Tx), receive from a node (Rx), or sleep, with the latter not requiring any specific cell selection [6]. The efficiency of the TSCH schedule is reflected in the consistent actions of devices. Thus, each device has to judiciously select the combination of (state, cell, neighbor), where the state$\in\left\{Tx,Rx,Sleep\right\}$. For instance, in node A’s TSCH schedule in Fig. 1, the cell (slot 1, channel 2) is set for receiving from node B. Conversely, in node B’s schedule, the same cell is allocated for transmitting to node A. Moreover, A is set to sleep mode during time slots 2 and 4, while B enters sleep mode in time slot 2.
@@
As the transition from IPv4 to IPv6 becomes more pressing due to IPv4’s address limitations and IPv6’s advanced capabilities, compatibility issues have risen. To tackle them, several bridging technologies have been developed to cope with the protocols’ distinct structures and behaviors. The latter became the de facto tools of the protocols’ coexistence, and thus ensuring a smooth transition from IPv4 towards IPv6. Within this framework, the Internet Engineering Task Force (IETF) took the initiative of forming the 6TiSCH working group, aiming to establish protocols that facilitate the implementation of IPv6 over the TSCH mode. The proposed protocol stack is depicted in Fig. 2. Specifically, the 6top layer holds the Scheduling Function (SF) responsible for TSCH cell allocation and traffic demands/performance metrics monitoring. The conventional SF is the Minimal Scheduling Function (MSF) [7]. Moreover, the Routing Protocol for Low-Power and Lossy Networks (RPL) layer, which encompasses the Objective Function (OF), plays a pivotal role in establishing and sustaining the routing topology. It does so by constructing a Destination-Oriented Directed Acyclic Graph (DODAG), a structured network topology that enables optimized routing. Through DODAG, RPL efficiently propagates routing information, tailoring pathways based on various critical metrics to meet the specific demands of the network. This mechanism is fundamental in addressing the complex routing challenges inherent in low-power and lossy networks, particularly within the Industrial IoT (IIoT) framework. The latter allows IoT nodes to discover their best potential parents and establish a route towards the network root. The conventional OF of RPL is the Objective Function Zero (OF0) [8].
Figure 2: The 6TiSCH protocol stack (Adapted from [9])
@@
Both RPL and 6top utilize control packets to maintain and optimize the IoT network’s operation, namely DODAG Information Solicitation (DIS), DODAG Information Object (DIO), and Destination Advertisement Object (DAO) for RPL, and 6P transactions in 6top. The relation between RPL and 6top is reciprocal. For instance, while RPL alters the routing topology by changing a node’s parent, 6top dynamically adjusts the scheduled cells to accommodate this change, i.e., releasing assigned cells to the old parent and negotiating new cell allocation with the new parent. RPL determines the next-hop parent for data transmission based on its topology, while 6top ensures cell availability for this specific transmission according to the TSCH schedule. However, this cross-layer interaction between the RPL and 6top layers (shown by red lines in Fig.2) controls implicitly very critical performance factors. To the best of our knowledge, very few works have investigated the nature of this interaction. Consequently, we propose a novel approach that improves the RPL-6top interaction, while aiming to reduce energy consumption and enhance the spatial reuse of resources. Our contributions can be summarized as follows:
  • We aim to reduce the number of control packets by merging two types of control packets and to simplify the cell negotiation and assignment process. Specifically, upon parent switching, RPL control messages will be used for cell allocation rather than transmitting separate 6P transaction messages.
  • We eliminate instances of parent switching failures, prevalent when the new parent schedule is at full capacity. This is achieved by embedding a list of available slots within the DODAG Information Object (DIO) packets. Consequently, a transition to a parent offering a superior rank will only be executed if the list of its free slots satisfies the requirement of the new child.
  • We improve the jitter and latency by adopting proximity-based slot selection rather than the conventional random selection.
  • To preempt packet drop events that occur when the TSCH queue reaches its limit, we propose setting a predefined queue size to solicit new cells from the parent node.
  • Through extensive simulations over the 6TiSCH simulator, we prove the superior performance of our proposed method in terms of joining time, energy consumption, traffic load, delivery ratio, and jitter, compared to the MSF benchmark.
The structure of this paper is outlined as follows: Section II presents a comprehensive overview of the research context and discusses related works. Section III delves into the intricate details of the proposed solution. Through simulations, section IV evaluates the performances of our approach and compares them to those of the MSF benchmark. Finally, Section V concludes the paper.
II Context and Related Work
@@
In the field of 6TiSCH networks, ongoing research is dedicated to developing effective, reliable, and energy-efficient communication strategies. This often involves examining and refining existing standards. In this section, we first outline the key conventional protocols within the 6TiSCH stack, thus providing the foundational framework for researchers. Then, we highlight recent related works.
II-A Context: 6TiSCH Background
@@
The 6TiSCH standards lay the groundwork for the operation of IIoT networks. We present their main components relevant to our research, namely the Minimal Configuration (MC) and the Minimal Scheduling Function (MSF). The MC details the procedure for the node joining process [10], while the MSF defines the scheduling mechanism, which outlines how communication cells are assigned and released [11].
II-A1 Minimal Configuration
@@
Figure 3: 6TiSCH network formation (MC).
  • Synchronization: Initially, each node is in a “pledge” state, passively listening on randomly selected channels to anticipate Enhanced Beacons (EBs) sent by nodes that are already part of the network, including the root node that initiate the network formation. Upon the reception of a predefined number (nb) of EB packets, a pledge synchronizes with the IoT network. For instance, node A in Fig. 3 starts as a “pledge” at time $T$ and synchronizes at time $T^{A}_{\rm Sync}$, where energy is expanded inefficiently during the interval $T_{1}=T^{A}_{\rm Sync}-T$.
@@
In summary, nodes within 6TiSCH networks dynamically alter their operational states throughout their life cycle, influenced by localized parameters and network conditions. These transitions and state changes, illustrated in Fig. 4, embody the node’s continual attempt to preserve network connectivity, synchronization, and data transmission integrity within the inherently challenging communication environment of 6TiSCH networks. Notably, a node reverts to its initial “pledge” state if it fails to receive any packet within a predefined temporal window (Dy_sync), indicative of potential communication disruptions or isolation from the network. Furthermore, if the node experiences a linkage failure to its parent or while executing parent switching (Dis_join), it regresses to a “synchronized” state, reflecting a lack of association with the DODAG, yet maintaining synchronization with the network’s time source.
@@
II-A2 Scheduling Function
The scheduling function, SF, serves as a critical mechanism to allocate time slots and channels within the TSCH mode. Its primary role is to manage the scheduling of communications across the time-frequency domain. By dynamically responding to the varying demands and conditions of the network, SF ensures the judicious use of resources and maintains dependable communications. As shown in Fig. 5, the SF conducts the cell reservation procedure using 6P protocol transactions [13, 11], where the term “Requester” denotes a node in need of reserving new cells from another node, referred to as the “Responder”. In this process, the “Requester” initiates the cell reservation process. There are two distinct approaches: the 2-step and 3-step methods. The key difference lies in which node proposes the list of cells. In the 2-step transaction illustrated in Fig. 4(a), the “Requester” proposes the list of cells, which leads to exchanging two control packets, whereas in the 3-step transaction, presented in Fig. 4(b), the “Responder” proposes the list of cells, thus exchanging three control packets. Nevertheless, in both approaches, the proposed cells are locked during the period $T_{\rm Lock}$. The conventional SF of the 6TiSCH protocol stack is the Minimal Scheduling Function, which follows a strategy of random cell selection, both in cell proposal during 6P transactions and in cell reservation among the proposed/available cells. Nevertheless, the nodes dynamically determine the number of cells needed based on their traffic requirements and network conditions. Although this stochastic approach provides adaptability to varying traffic conditions, it may face challenges such as internal collisions, particularly in dense networks. This raises concerns about its applicability in scenarios with stringent Quality-of-Service (QoS) requirements.
(a) 2-Step 6P Transaction.

### v2

@@ Abstract
The 6TiSCH protocol stack plays a vital role in enabling reliable and energy-efficient communications for the Industrial Internet of Things (IIoT). However, it faces challenges, including prolonged network formation, inefficient parent switching, high control packet overhead, and suboptimal resource utilization. To tackle these issues, we propose in this paper a novel cross-layer optimization framework aiming to enhance the coordination between the Scheduling Function (SF), the Routing Protocol for Low-Power and Lossy Networks (RPL), and queue management. Our solution introduces a slot-aware parent switching mechanism, early slot reservation to mitigate queue overflow, and a refined slot locking strategy to improve slot availability. To reduce control overhead, the proposed method merges 6P cell reservation information into RPL control packets (DIO/DAO), thus minimizing control exchanges during parent switching and node joining. Optimized slot selection further reduces latency and jitter. Through extensive simulations on the 6TiSCH simulator and under varying network densities and traffic loads, we demonstrate significant improvements over the standard 6TiSCH benchmark in terms of traffic load, joining time, latency, and energy efficiency. These enhancements make the proposed solution suitable for time-sensitive IIoT applications.
keywords:
Industrial IoT, TSCH, 6TiSCH, SF, RPL, TSCH Queue Management, Cross-Layer Optimization, Network Formation, Energy Efficiency.
[…]
[inst1]organization=University of Biskra, Dept. Computer Science,addressline=BP 145 RP, city=Biskra, postcode=07000, state=Biskra, country=Algeria
[…]
[inst2]organization=École de Technologie Supérieure (ETS), Dept. Software and IT Engineering,addressline=1100 R. Notre Dame W., city=Montreal, postcode=H3C 1K3, state=Quebec, country=Canada
[…]
[inst4]organization=University of Batna, Dept. Computer Science,addressline=Univ. Batna, city=Batna, postcode=05000, state=, country=Algeria
1 Introduction
The Industrial Internet of Things (IIoT) is reshaping critical sectors such as industrial automation, smart grids, and environmental monitoring by enabling large-scale, real-time data exchange across vast sensor networks. However, achieving ultra-reliable, low-latency, and energy-efficient communication in such resource-constrained environments remains a major challenge. Current networking protocols struggle to provide deterministic performance while balancing energy efficiency, scalability, and real-time constraints, making protocol enhancements crucial for IIoT applications. To meet these stringent requirements, 6TiSCH has emerged as a key protocol stack, integrating IPv6 over the Time-Slotted Channel Hopping (TSCH) mode of IEEE 802.15.4e. By leveraging time-slotted access, multi-channel communication, and channel hopping, 6TiSCH ensures deterministic reliability, enhances interference mitigation, and provides efficient energy management, making it a strong candidate for time-sensitive IIoT applications.
The 6TiSCH protocol encapsulates IPv6 over the TSCH mode of IEEE 802.15.4e, marking a significant advancement in networking infrastructure for IIoT. This protocol merges the precise and predictable TSCH mechanism with the broad capabilities of IPv6, thus providing industrial-level reliability and precise timing. Indeed, 6TiSCH is designed for low-power and lossy networks (LLNs), which are typical in IIoT environments due to their limited power and unreliable transmission media. This makes it particularly effective for seamless integration and interoperability among low-power IoT devices across various environments.
6TiSCH ensures reliable, low-latency communication, making it particularly suitable for time-sensitive applications such as industrial automation, healthcare, and environmental monitoring. Its structured scheduling mechanism supports real-time data transmission and synchronization, which are critical for these domains. However, existing scheduling approaches lack adaptability, leading to inefficient slot utilization and degraded real-time performance in IIoT deployments. Indeed, for real-time IIoT applications, minimizing jitter and latency is essential to ensure stable packet transmission and synchronization. However, inefficiencies in 6TiSCH’s scheduling and queue management mechanisms introduce unpredictable delays, negatively affecting control loops and time-sensitive decision-making processes. Moreover, energy efficiency in IIoT stands at the forefront of technological challenges, with researchers globally striving to meld enhanced performances with reduced energy consumption. Specifically, the medium access control (MAC) layer is pivotal in controlling the energy drain by regulating access to the shared communication medium, typically a wireless channel. The TSCH protocol stands out as the leading MAC operational mechanism and is included in several standards, such as ANSI/ISA100.11a and WirelessHart IEC62591 [1], [2].
@@ 1 Introduction
The core concept of TSCH lies in integrating Time Division Multiple Access (TDMA) with channel-hopping techniques. This fusion results in a dual medium access mechanism based on time and channel. In this context, “Time” is referred to by time slot, and “Frequency” is referred to by channel. A time slot within a specific channel is defined as a cell. For every time slot, shown in Fig. 1, an IoT node specifies its activity in a given channel, choosing to transmit to a node (Tx), receive from a node (Rx), or sleep, with the latter not requiring any specific cell selection [3]. Moreover, a “slotframe” represents a repeating schedule of time slots. The efficiency of the TSCH schedule is reflected in the consistent actions of devices. Thus, each device has to judiciously select the combination of (state, cell, neighbor), where the state $\in\left\{Tx,Rx,Sleep\right\}$. For instance, in node $A$’s TSCH schedule in Fig. 1, the cell (slot 1, channel 2) is set for receiving from node $B$. Conversely, in node $B$’s schedule, the same cell is allocated for transmitting to node $A$. Moreover, $A$ is set to sleep mode during time slots $2$ and $4$, while $B$ enters sleep mode in time slot $2$.
Figure 2: The 6TiSCH protocol stack [4].
@@
Building on TSCH, the Internet Engineering Task Force (IETF) established the 6TiSCH working group to define protocols that seamlessly integrate IPv6 over TSCH, enabling efficient network formation, routing, and resource scheduling in low-power industrial environments. The 6TiSCH protocol stack is depicted in Fig.2. Specifically, the 6top layer holds the Scheduling Function (SF) responsible for TSCH cell allocation and traffic demands/performance metrics monitoring. The conventional SF is the Minimal Scheduling Function (MSF) [5]. Moreover, the Routing Protocol for Low-Power and Lossy Networks (RPL) layer, which encompasses the Objective Function (OF), establishes and sustains the routing topology. It does so by constructing a Destination-Oriented Directed Acyclic Graph (DODAG), a structured network topology that enables optimized routing. Through DODAG, RPL efficiently propagates routing information, tailoring pathways based on various critical metrics to meet the specific demands of the network. Also, RPL allows IoT nodes, in particular Industrial IoT (IIoT), to discover their best potential parents and establish a route towards the network root. The conventional OF of RPL is the Objective Function Zero (OF0) [6].
@@
Both RPL and 6top utilize control packets to maintain and optimize the IoT network’s operation, namely DODAG Information Solicitation (DIS), DODAG Information Object (DIO), and Destination Advertisement Object (DAO) for RPL, and 6P transactions in 6top. The relation between RPL and 6top is reciprocal. For instance, while RPL alters the routing topology by changing a node’s parent, 6top dynamically adjusts the scheduled cells to accommodate this change, i.e., releasing assigned cells to the old parent and negotiating new cell allocation with the new parent. Moreover, RPL determines the next-hop parent for data transmission based on its topology, while 6top ensures cell availability for this specific transmission according to the TSCH schedule. However, this cross-layer interaction between the RPL and 6top layers (shown by red lines in Fig.2) implicitly controls very critical performance metrics, such as the joining time of a new node to the DODAG network and energy efficiency.
At the intersection of SF and RPL, inefficiencies in routing and scheduling coordination result in wasted network resources, increased energy consumption, and higher packet loss. A major contributor to this problem is the inefficient parent switching mechanism. In 6TiSCH, nodes may switch to a new parent without verifying slot availability, leading to network instability and inefficient resource allocation. Another challenge arises when a node joins the network but experiences significant delays before becoming fully operational. Upon joining the DODAG, a node must undergo multiple control packet exchanges before it can transmit data. Since these packets are sent in separate time slots, the process of becoming operational is delayed, preventing the node from contributing to network traffic immediately. This inefficiency is exacerbated in dense networks, where multiple joining nodes compete for minimal shared cells, further extending the delay. Beyond SF-RPL coordination challenges, inefficiencies in the coordination between queue management and slot allocation can degrade network performance. A major issue is queue overflow, which occurs when a node accumulates packets but lacks available transmission cells, leading to packet drop. Specifically, when the transmission cells to the destination are lacking, the queue fills up, causing packet drops before transmission. Finally, a critical inefficiency stems from slot reservation delays caused by cell locking. Specifically, when a node proposes a set of cells for allocation, the latter are locked until either receiving a response or timing out. During this period, the slots remain unavailable for reservation, leading to resource losses and delayed slot reassignment.
To address these challenges, we propose a novel cross-layer optimization framework designed to enhance 6TiSCH network performances. To the best of our knowledge, very few works have investigated the cross-layer design of scheduling (6top) and routing (RPL) functions within 6TiSCH networks and their generated advantages and limitations. Hence, the main contributions of this work can be summarized as follows:
  1. 1.
    We propose a novel slot-aware parent switching mechanism where nodes check slot availability before switching, ensuring more stable routing decisions and reducing unnecessary reconfigurations.
  2. 2.
    We introduce parallel control packet transmissions to reduce the node activation time.
  3. 3.
    To attenuate packet drops, we develop an early slot reservation strategy where nodes dynamically request additional cells before queue overflow occurs.
  4. 4.
    To improve slot availability and resource usage, we propose a flexible slot locking mechanism that limits the duration for which proposed slots remain reserved without confirmation.
  5. 5.
    Through extensive simulations in the 6TiSCH simulator [7], we demonstrate the efficacy of our framework in improving the joining time, energy efficiency, packet delivery, and jitter performance, making it a robust enhancement for IIoT.
The remainder of this paper is structured as follows: Section 2 provides background on 6TiSCH and reviews related works. Section 3 identifies key challenges and inefficiencies in 6TiSCH networks. Section 4 details the proposed solution, explaining its design and implementation. Section 5 evaluates the performance of our approach through simulations, comparing it to the MSF benchmark. Section 6 discusses the challenges and future research directions. Finally, Section 7 concludes the paper.
2 6TiSCH Background and Related Works
@@
In the field of 6TiSCH networks, ongoing research is dedicated to developing effective, reliable, and energy-efficient communication strategies. This often involves examining and refining existing standards. In this section, we first outline the key conventional protocols within the 6TiSCH stack.Then, we highlight recent related works.
2.1 6TiSCH Background
@@
The 6TiSCH standard lays the groundwork for the operation of IoT networks. We present their main components relevant to our research, namely the Minimal Configuration (MC) and the Minimal Scheduling Function (MSF). The MC details the procedure for the node joining process [8], while the MSF defines the scheduling mechanism, which outlines how communication cells are assigned and released [9].
2.1.1 Minimal Configuration (MC)
@@
Figure 3: 6TiSCH network formation using MC.
  • Synchronization: Initially, each node is in a “pledge” state, passively listening on randomly selected channels to anticipate Enhanced Beacons (EBs) sent by nodes that are already part of the network, including the root node that initiates the network formation. Upon the reception of a predefined number of EB packets, a “pledge” synchronizes with the IoT network. For instance, node A in Fig. 3 starts in the “pledge” status at time $T$ and synchronizes at time $T^{A}_{\rm Sync}$, where energy is expanded during the interval $T_{1}=T^{A}_{\rm Sync}-T$.
@@
In summary, nodes within 6TiSCH networks dynamically alter their operational states throughout their life cycle, influenced by localized parameters and network conditions. These transitions and state changes, illustrated in Fig. 4, embody the node’s continual attempt to preserve network connectivity, synchronization, and data transmission integrity within the inherently challenging communication environment of 6TiSCH networks. Notably, a node reverts to its initial “pledge” state if it fails to receive any packet within a predefined temporal window (Dy_sync), indicative of potential communication disruptions or isolation from the network. Furthermore, if the node experiences a linkage failure to its parent or while executing parent switching (Dis_join), it regresses to a “synchronized” state.
@@
2.1.2 Scheduling Function (SF)
The scheduling function, SF, serves as a critical mechanism to allocate time slots and channels (i.e., cells) within the TSCH mode. Its primary role is to manage the scheduling of communications across the time-frequency domain. By dynamically responding to the varying demands and conditions of the network, SF ensures the judicious use of resources and maintains dependable communications. As shown in Fig. 5, the SF conducts the cell reservation procedure using 6P protocol transactions, where the term “Requester” denotes a node in need of reserving new cells from another node, referred to as the “Responder”. In this process, the “Requester” initiates the cell reservation process [11, 9]. There are two distinct approaches: the 2-step and 3-step methods. The key difference lies in which node proposes the list of cells. In the 2-step transaction illustrated in Fig. 5(a), the “Requester” proposes the list of cells, which leads to exchanging two control packets, whereas in the 3-step transaction, presented in Fig. 5(b), the “Responder” proposes the list of cells, thus exchanging three control packets. Nevertheless, in both approaches, the proposed cells are locked during the period $T_{\rm Lock}$. The conventional SF of the 6TiSCH protocol stack is the Minimal Scheduling Function (MSF), which follows a strategy of random cell selection, both in cell proposal during 6P transactions and in cell reservation among the proposed/available cells. Nevertheless, the nodes dynamically determine the number of cells needed based on their traffic requirements and network conditions. Although this stochastic approach provides adaptability to varying traffic conditions, it may face challenges such as internal collisions, particularly in dense networks. This raises concerns about its applicability in scenarios with stringent Quality-of-Service (QoS) requirements.
(a) 2-Step 6P transaction